AO Commissioning at RFBG: 26 November 2010 UTC

Observer(s): SEsposito, APuglisi, LBusoni, EPinna, DMiller, GBrusa (LBTO)

Telescope Operator: DHuerta (LBTO)

SW Support: TLiebold (Tucson)

Telescope Support: JLittle (LBTO)

Instrument Support: JMorris (LBTO)

Summary

Preparations

Last night AO guys turned one of the coolant values off inside the W-Unit that cools the Pupil Positioning Lens electronics because it was too cold and the lens was not moving. It is working fine so far this evening.

20:30 John filled LBC and MODS

Details

00:30 David opened the Chamber

00:40 Set IE=-24, CA=41 Collimate by hand a bit

00:45 Collimate on a BS star

01:00 newpoint AO794 ACE GS=6 PA=90 (oops, Dave sent 792. Send again 794) No guide star found

01:03 preset AO794 GS=0 ACQUIRE set IE=-34 CA

01:05 preset AO792 ACE GS=6 PA=90 Failed, AO xy positions were in the millions (again, not the star we wanted)

01:07 preset AO794 ACE GS=6 PA=90 Failed, AO xy positions were in the millions

IRAF Logs: /home/telescope/iraf/LOG/newpoint_010851.log, Misc/target_010851.dat, Misc/reference_010851.dat

newpoint_010851.log
   LBTtools -->>  Observe:newpoint.cl  <<-- Fri 01:08:51 26-Nov-2010

* Current working directory: /home/telescope/iraf
* Time since 2000.0 in years: 10.90183
* CatDir = /lbt/catalogs
* Found catalog: /AO/catalogs/AO794.catalog

Science target: AO794
   22:40:52.4720 20:49:22.210  6.20 -0.40  2000.0 J2000  7.38  2.25  
   
Guide star: GS06
   22:41:00.9170 20:50:15.370  0.00 0.00  2000.0 J2000  13.93  1.39  
   Offset: AO794->GS06 = 129.72 arcsec (118.33 E 53.16 N)
   Use PA=65.81 for AGw theta=0

AO reference star: AOref*00
   22:40:52.4720 20:49:22.210  6.20 -0.40  2000.0 J2000  7.38  2.25  
   Offset: AO794->AOref*00 = 0.00 arcsec (0.00 W 0.00 S)

   #################################################################
   ###  Sending the following command sequence through the IRC:  ###
   #################################################################
   IRCR ClearHotspot

   IRCR ClearOffset

   IRCR ClearStars > /dev/null ; echo $?
   IRCR ClearStars
   IRCR SetStars Misc/target_010851.dat

   IRCR ClearReference > /dev/null ; echo $?
   IRCR ClearReference
   IRCR SetReference Misc/aorefstar_010851.dat

   IRCR PresetTelescope 90. POSITION ADAPTIVEACE right

   IRCR RunAO false right


...Done!
   to examine the log: page LOGs/newpoint_010851.log
target_010851.dat
# Target coordinates for source AO794
22:40:52.4720 20:49:22.210  RADEC 2000. J2000 6.2 -0.4 7.38 R 2.25 B_V 1.4
# Guide Star coordinates for GS06
22:41:00.9170 20:50:15.370  RADEC 2000. J2000 0. 0. 13.93 R 1.39 B_V 0.63
aorefstar_010851.dat
# AO Reference Star coordinates for AOref*00
22:40:52.4720 20:49:22.210  RADEC 2000. J2000 6.2 -0.4 7.38 R 2.25 B_V 0.63

01:10 From IDL Doug sends: preset, AO794, ACE, GS=6, RS=0, PA=90 Failed: Framerate 4 is out of range....bla bla bla

PCS_XY values sent to the AOS were correct

IDL Output
IDL> iif_preset, AO794, MODE=ACE, GS=6, RS=0, PA=90
  
      # Target star: 0   AO794
        22:40:52.4787 +20:49:22.206 RADEC 2000.0 J2000 0.0 0.0 7.38 U 2.25 B_V 0.3
      # Guide star: 6
        22:41:00.9170 +20:50:15.370 RADEC 2000.0 J2000 0.0 0.0 13.93 U 1.39 B_V 0.5
      # AO Reference star: 0
        22:40:52.4787 +20:49:22.206 RADEC 2000.0 J2000 0.0 0.0 7.38 U 2.25 B_V 0.5
  
      # Y-Axis Position Angle:   65.79
  
        request = irc PresetTelescope 90.00 POSITION ADAPTIVEACE right; echo $?
IDL> 

01:15 Alfio tweak parameters and close AO loop. Seeing is bad, 1.5-2.5" fwhm on guider

02:17 Seeing is still bad so decide to test presets in ADAPTIVE mode at low elevations: 50,40,30

02:19 preset, AO704, RS=0, GS=0,

02:27:22 preset, AO704, RS=0, GS=0, MODE=ACE Failed because star too dim: Actually the star is 4" long (astigmatism) and some fo the light is blocked by the pinhole in front of the Pyamid.

02:28 AOS died as well as all the Adaptive Secondary processes.

02:40:53 preset, AO704, RS=0, GS=0, MODE=ACE Failed because wfs system was not ready

02:42:13 preset, AO704, RS=0, GS=0, MODE=ACE Success 02:43:32

02:43:32 Collimate 4 times

02:46:52 irc_runao Failed

02:50:46 preset, AO704, RS=0, GS=3, MODE=ACE Success 02:51:30

02:56:36 irc_runao Failed: guide star too dim

02:59:10 preset, AO767, RS=0, GS=3, MODE=ACE Success 02:59:08 Elevation = 47

03:03:35 irc_runao Failed and guide star moved out of the GCS guide window. I turned on guiding pulled star back. Alfio is now working on AO engineering GUI's

03:06 Alfio found that the AO supervisor was not getting the rotator angle. GCS does have the correct angle. The bug is searched.

03:50 Found and fixed the rotator software problem

03:52:16 preset, AO767, RS=0, GS=3, MODE=ACE Success 03:52:51

03:53 Collimate a few times

03:58:30 irc_runao 04:00:45 Success, but loop opend immediately

04:01:22 preset, AO767, RS=0, GS=3, MODE=ACE Success 04:02:00

04:02:17 irc_runao Success 04:05:00 Loop opened a few seconds later

04:43 Alfio found that the intensity measure by the Pyamid was changing by a factor of two because of 1.5 to 3" seeing. Could not keep the loop closed on a 7.0 star at 30 elevation in bin1 mode. Changed to bin2 and close loop was stable.

04:45 Collect IRCT data

*04:

05:05:29 preset, AO74, RS=0, GS=1, MODE=ACE, PA=90 Success 05:06:44

05:11:32 irc_runao Success 05:14:25 Guider 1-1.4" fwhm

07:56 preset, AO1076, ACQUIRE, GS=0

07:57:40 preset, AO1076, ACE, RS=0, GS=4 Success 07:58:55

08:00:05 irc_runao Success 08:02:32

08:40 Alfio's data collection script stop and IRTC GUI has a lib error. He fixes the bug and restarts the IRTC

08:42 David preset in ACQUIRE mode and sets IE=-53 CA=38

08:44 preset , AO1076, ACE, RS=0, GS=4 Success

08:45 irc_runao Failed no ref star found

08:46 preset , AO1076, ACE, RS=0, GS=4 Success

08:47 Alfio tweaks some parameters

08:48 preset , AO1076, ACE, RS=0, GS=4 Success

08:49 irc_runao

08:50 AO guys collect a bunch of data

09:55 David puts the telescope to horizon so Neutral Density can be place in front of W-Unit

10:20 Neutral Density is in. David set IE=-65 CA=20

10:38 Collimated after 10 iterations. We seemed to have a bunch of coma. However, we were collimated at 60 elevation before we moved to horizon. After we preset to 75 elevation we had big coma. Why? David cleared C00 on the primary, but no change.

10:40 irc_runao Success

11:18 preset , AO1076, ACE, RS=0, GS=4 Success

11:20 irc_runao

12:35 preset , AO327, ACE, RS=0, GS=7 Success

12:40 irc_runao

13:00 Closed

-- DougMiller - 26 Nov 2010
Topic revision: r8 - 27 Nov 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback